OpenStudy (victoriaf16):

Shameka is tossing acorns into a hole. She started with 800 acorns in the hole. On average, she gets 75 acorns per hour in the hole. Write an equation that represents the amount of acorns, A, in the hole, where h is the number of hours. This is an awful question that I'm having trouble with, please help!

OpenStudy (kropot72):

When she starts throwing acorns into the hole, there are already 800 acorns in the hole. So one hour after she starts throwing, how many acorns would be in the hole?

OpenStudy (victoriaf16):

oops I left out the answer choices they are; A=75h-800 A=75h+800 A=-800h-75 A=-800h+75